### v1

# and non-Abelian anyons

Abstract
We introduce a frequency–tunable, two–dimensional non-Abelian control over operation order built from the reduced Burau representation of the braid group $B_{3}$, specialised at $t=e^{i\omega}$ and unitarized by Squier’s Hermitian form. Coupled to two non-commuting qubit unitaries $A,B$, the resulting switch admits a closed expression for the single-shot Helstrom success and a fixed-order ceiling $p_{\rm fixed}^{*}$, yielding an explicit, analytic witness gap $\Delta(\omega)=p_{\rm switch}(\omega)-p_{\rm fixed}^{*}$. We prove that $\Delta(\omega)>0$ is achievable, thereby certifying causal non-separability by purely algebraic means, and confirm this behaviour numerically. Conceptually, this furnishes a minimal non-Abelian $B_{3}$ control for a Gedankenexperiment in anyonic statistics.
1 Introduction
Indefinite causal order (ICO) is typically modeled by the quantum switch- a coherent superposition of two operation orders, $AB$ and $BA$ [1, 2]. Most certifications of ICO employ the process-matrix framework and causal witnesses formulated as semidefinite programs (SDPs) [3, 4].
While there are frameworks [4, 5, 6, 7] for indefinite causal order that involve three observers – commonly known as Alice, Bob, and Charlie – only two operations are coherently superposed in order. Thus, even though three parties appear, the algebraic structure remains that of a single “swap-order” generator, effectively equivalent to the braid group $B_{2}\simeq\mathbb{Z}$. By contrast, the minimal non–Abelian braid group $B_{3}$ has two independent generators $\sigma_{1},\sigma_{2}$ obeying the Yang–Baxter relation $\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}$, and admits non–trivial two–dimensional unitary representations that are matrix– rather than phase–valued, marking the transition from Abelian order interference to genuine non–Abelian braiding.
Recent theoretical and experimental works have begun to explore braiding phenomena in non-Abelian or non-Hermitian contexts, including three-band acoustic lattices [8], non-Abelian photonic braid monopoles [9], and programmable photonic lattices for non-Abelian topology [10]. These studies realize eigenvalue braids or non-unitary transfer operators in parameter space, but they do not address the explicit $B_{3}$ representation acting unitarily on a two-dimensional control space.
Here we introduce a representation-theoretic model of coherent order control based on the reduced Burau representation $\psi(w)$ of $B_{3}$, specialized at $t=e^{i\omega}$ and rendered unitary by Squier’s Hermitian form [11, 12].
Within this framework, the control mixer $M(\omega)$ is a $2\times 2$ unitary obtained from a braid word $w$, and the overall switch combines two non-commuting target operations $A,B\in U(2)$ through a controlled superposition of orders. The resulting process $T(\omega)$ admits closed-form discrimination probabilities through Helstrom’s theorem [13], allowing us to define an analytic witness gap $\Delta(\omega)$ that certifies causal non-separability when positive.
Unlike $B_{2}$-based or non-Hermitian constructions, the present formulation yields a minimal, fully unitary $B_{3}$–based Gedankenexperiment in which non-commuting braid generators couple directly to non-commuting target operations. This establishes a concrete mathematical route by which non-Abelian exchange statistics could be detected through order-sensitive interference.
2 Manuscript contents
Section 3 sets up the Burau–Squier control and the switch. Section 6 defines the task and discusses the Helstrom formula. Section 7 contains the main theorem: closed-form witness constructed from the braid group, while the reproducible code is available on GitHub [14].
3 Braid group control and the switch
[…]
Braid group on three strands.
Let $B_{3}$ denote the braid group on $3$ strands. There are several equivalent definitions of this mathematical object.
Topological / combinatorial. $B_{3}$ is the group of ambient isotopy classes of geometric 33–strand braids in the cylinder $D^{2}\times[0,1]$, with group law given by vertical concatenation (stacking). Cf. [15, 16].
Abstract group presentation. $B_{3}$ is generated by elementary crossings $\sigma_{1},\sigma_{2}$ subject to the Yang–Baxter relation
$B_{3}\;\cong\;\big\langle\sigma_{1},\sigma_{2}\ \big|\ \sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}\big\rangle.$ (1)
Mapping class group. $B_{3}\cong\operatorname{MCG}(D^{2}\setminus\{3\text{ marked points}\})$, the mapping class group of a thrice-punctured disc where braids are represented as motions of punctures (or mutually distinguishable marked point) with time [15].
Pictures.
The generators and a sample word are depicted in Figure 1: black curves are strands; time flows upward. We place strand 1 on the left, 2 in the middle, and 3 on the right. Thus, the picture of $\sigma_{1}$, resp. $\sigma_{2}$, is the positive crossing of strands $(1,2)$, resp. $(2,3)$.
$\sigma_{1}$$\sigma_{2}$$w=\sigma_{1}\,\sigma_{2}\,\sigma_{1}$
Figure 1: Two braid generators of $B_{3}$, $\sigma_{1}$ and $\sigma_{2}$, and their product $w=\sigma_{1}\sigma_{2}\sigma_{1}$ (which satisfies $w=\sigma_{2}\sigma_{1}\sigma_{2}$, up to strand rearrangement).
Reduced Burau representation for $B_{3}$.
For $w\in B_{3}$ and a parameter $t\in\mathbb{C}^{\times}$, the reduced Burau representation
$\psi:\ B_{3}\longrightarrow\mathrm{GL}\!\big(2,\mathbb{Z}[t,t^{-1}]\big)$
is defined on generators by
$\psi(\sigma_{1})\;=\;\begin{pmatrix}-t&1\\[2.0pt] 0&1\end{pmatrix},\qquad\psi(\sigma_{2})\;=\;\begin{pmatrix}1&0\\[2.0pt] t&-t\end{pmatrix}$, (2)
and extended by way of homomorphism:
$\psi(\sigma_{i_{1}}\sigma_{i_{2}}\ldots\sigma_{i_{k}})=\psi(\sigma_{i_{1}})\,\psi(\sigma_{i_{2}})\,\ldots\,\psi(\sigma_{i_{k}}).$ (3)
The generators given above are the standard $2\times 2$ reduced Burau matrices; see [15, § 3] or [16, § 3.1].
Faithfulness for $B_{3}$.
The reduced Burau representation can be defined for any $n\geq 2$, and is known to be faithful (i.e. a group isomorphism) for $n\leq 3$ [16, § 3.3.2]; by contrast, it is not faithful for $n\geq 5$ [17, 18, 19]. For $n=4$, it is still an open problem. In general, braid groups are known to be linear, i.e. they admit a faithful linear representation [20, 21]. However, this representation does not have its image inside $\mathrm{GL}\!(2)$.
4 Squier’s unitarization
Let $s$ be an indeterminate with involution $\overline{s}=s^{-1}$ and set $t=s^{2}$. Squier’s modification of the reduced Burau representation is the homomorphism
$\beta:\ B_{3}\longrightarrow\mathrm{GL}\!\big(2,\mathbb{Z}[s,s^{-1}]\big),\qquad\sigma_{i}\longmapsto\beta_{i}=\beta(\sigma_{i}),$
defined on generators by
$\beta_{1}=\beta(\sigma_{1})\;=\;\begin{pmatrix}-s^{2}&s\\[2.0pt] 0&1\end{pmatrix},\qquad\beta_{2}=\beta(\sigma_{2})\;=\;\begin{pmatrix}1&0\\[2.0pt] s&-s^{2}\end{pmatrix}$. (4)
On the coefficient ring $\mathbb{Z}[s,s^{-1}]$ we use the *–involution given by entry-wise $\overline{\cdot}$ and transpose: $M^{\ast}:=\overline{M}^{\,T}$, with $\overline{s}=s^{-1}$.
Squier exhibits the Hermitian form
$J(s)\;=\;(s+s^{-1})\,I_{2}\;-\;\begin{pmatrix}0&1\\[2.0pt] 1&0\end{pmatrix}\;=\;\begin{pmatrix}s+s^{-1}&-1\\[2.0pt] -1&s+s^{-1}\end{pmatrix}$, (5)
and proves the identities
$\beta_{i}(s)^{\ast}\,J(s)\,\beta_{i}(s)\;=\;J(s)\qquad(i=1,2).$ (6)
After specialization $s=e^{i\omega/2}$, and hence $t=s^{2}=e^{i\omega}$, equation (6) gives us ordinary $J(\omega)$–unitarity.
Relation to the standard reduced Burau.
Let us set
$D(s)=\mathrm{diag}(1,s^{-1}).$
Then
$\psi(\sigma_{i})\;=\;D(s)^{-1}\,\beta(\sigma_{i})\,D(s),\qquad\tilde{J}(s)\;=\;D(s)^{\ast}\,J(s)\,D(s),$ (7)
and consequently $\psi(\sigma_{i})^{\ast}\,\tilde{J}(s)\,\psi(\sigma_{i})=\tilde{J}(s)$.
Unitary specializations and the control parameter.
Given a phase $\omega\in[0,2\pi]$, we specialize the representation at $s=e^{i\omega/2}$. The Hermitian form defined in (5) becomes
$J(\omega)\;=\;\begin{pmatrix}2\cos(\omega/2)&-1\\[2.0pt] -1&2\cos(\omega/2)\end{pmatrix}$. (8)
Its eigenvalues are $\lambda_{\pm}(\omega)=2\cos(\omega/2)\mp 1,$ hence $J(\omega)\succ 0$ precisely for
$|\omega|\;<\;\tfrac{2\pi}{3}\quad(\mathrm{mod}\ 2\pi).$ (9)
On any open interval where $J(\omega)\succ 0$, compute $J(\omega)=R(\omega)^{\dagger}R(\omega)$ using the Cholesky decomposition. For any braid word $w\in B_{3}$ we then obtain an ordinary Euclidean unitary by means of matrix similarity
$U(\omega)\;=\;R(\omega)\,\beta(w)\,R(\omega)^{-1}\ \in\ U(2),$ (10)
since (6) implies $U(\omega)^{\dagger}U(\omega)=I_{2}$. We henceforth restrict to nontrivial specializations, i.e. $\omega$ in the positivity window (9) and braid words $w$ for which $\beta(w)$ does not collapse to a scalar multiple of $I_{2}$.
5 Switch block and test device
Given $\omega$ such $J(\omega)\succ 0$, and $J=R^{\dagger}R$, set the mixer
$M(\omega)\ :=\ R(\omega)\,\beta(w)\big|_{s=e^{i\omega/2}}\,R^{-1}(\omega)\ \in U(2).$ (11)
Define the switch as
[…]
where $\theta=\theta(\omega)$ is an arbitrary phase map.
The full test device is
$T(\omega)\ =\ \big(M_{\rm post}(\omega)\otimes I\big)\ S(\theta(\omega))\ \big(M_{\rm pre}(\omega)\otimes I\big),$ (13)
where we allow distinct braid words to produce $M_{\rm pre}(\omega)$ and $M_{\rm post}(\omega)$.
Lemma 5.1 (Squier positivity intervals).
For any fixed braid words $w_{\mathrm{pre}/\mathrm{post}}\in B_{3}$, the set
$\Omega_{+}\ :=\ \{\omega\in[0,2\pi):J(\omega)\succ 0\}$
is a finite union of open intervals whose endpoints lie in $Z$. On each connected component $I\subset\Omega_{+}$ one can choose $R(\omega)$ with $J(\omega)=R(\omega)^{\dagger}R(\omega)$ that depends real-analytically on $\omega$, and consequently
$M_{\mathrm{pre}}(\omega)=R(\omega)\,\beta_{3}(w_{\mathrm{pre}})\,R(\omega)^{-1},\quad M_{\mathrm{post}}(\omega)=R(\omega)\,\beta_{3}(w_{\mathrm{post}})\,R(\omega)^{-1}$
are U(2)U(2)–valued, real-analytic functions of $\omega$ on $I$.
Proof.
$J(\omega)$ has real-analytic entries in $\omega$, hence $\det J(\omega)$ is a real-analytic $2\pi$–periodic function; its zero set $Z$ is finite unless $\det J\equiv 0$ (which is not the case here). Thus $\Omega_{+}$ is a finite union of open intervals with endpoints in $Z$. On a compact subinterval of $I$, the spectrum of $J(\omega)$ is uniformly bounded away from 0, so the Cholesky factor $R(\omega)$ may be chosen to be real-analytic, cf. [22, Chapter II, §6.1–6.4]. Conjugation by such a matrix $R(\omega)$ preserves analyticity, and JJ–unitarity of $\beta_{3}(\cdot)$ implies $M_{\mathrm{pre/post}}(\omega)\in U(2)$. ∎
Proposition 5.2 (Smoothness of the whole device).
On every positivity interval $I\subset\Omega_{+}$ from Lemma 5.1, the map $\omega\mapsto T(\omega)$ is real-analytic as a U(2)U(2)–valued function. In particular, all matrix entries of $T(\omega)$, its eigenvalues and eigenphases, spectral projectors, and any polynomial or rational functions thereof are real-analytic in $\omega$.
Proof.
By Lemma 5.1, $M_{\mathrm{pre/post}}(\omega)$ are real-analytic on $I$. The map $\omega\mapsto e^{i\theta(\omega)}$ is smooth and real-analytic, if $\theta$ is. Hence, so is $S(\theta(\omega))$. Products of real-analytic matrix-valued functions are real-analytic. ∎
6 Helstrom’s formula and convexity witness
In order to determine if more than just interference of two causal orders takes place in the test device, we use single-shot binary discrimination with equal priors.
Helstrom for unitaries.
Let $U_{0},U_{1}$ be two unitaries. The optimal single-shot success is
$p^{*}(U_{0},U_{1})\ =\ \frac{1}{2}\Big(1+\sin\!\big(\tfrac{1}{2}\Delta\big)\Big),$ (14)
where $\Delta\in[0,\pi]$ is the shortest eigenphase arc (spectral spread) of $U_{0}^{\dagger}U_{1}$. Equation (14) follows from the Helstrom bound [13] specialized to two unitaries $U_{0}$, $U_{1}$ with equal priors.
Lemma 6.1 (Helstrom formula for unitaries).
Let $U_{0},U_{1}\in U(2)$ and define $V=U_{0}^{\dagger}U_{1}$. Let $\Delta\in[0,\pi]$ be the length of the smallest closed arc on the unit circle containing the spectrum of $V$. Then the optimal single-shot, equal-prior success probability for distinguishing $U_{0}$ and $U_{1}$ is
$p^{*}(U_{0},U_{1})=\tfrac{1}{2}\Big(1+\sin(\Delta/2)\Big).$
Proof.
[…]
For a rank-1 projector, the trace norm is computable in closed form:
$\bigl\|\,|\psi\rangle\!\langle\psi|-V|\psi\rangle\!\langle\psi|V^{\dagger}\,\bigr\|_{1}=2\sqrt{\,1-\big|\langle\psi|V|\psi\rangle\big|^{2}\,}.$
Hence
$p^{*}(U_{0},U_{1})=\tfrac{1}{2}\!\left(1+\sqrt{\,1-\rho(V)^{2}\,}\right),\qquad\rho(V):=\min_{|\psi\rangle}\!\left|\langle\psi|V|\psi\rangle\right|.$
Because $V$ is unitarily diagonalisable, its numerical range $W(V)=\{\langle\psi|V|\psi\rangle:\|\psi\|=1\}$ is the convex hull of its eigenvalues: all $\{e^{i\theta_{k}}\}$ are situated on the unit circle. The minimal modulus $\rho(V)$ is therefore the Euclidean distance from the origin to $W(V)$.
If the eigenvalues of $V$ occupy a shortest covering arc of angular width $\Delta<\pi$, then $W(V)$ is a circular arc’s convex hull, and the closest point to the origin is the midpoint of the chord joining $e^{\pm i\Delta/2}$, giving
$\rho(V)=\cos(\Delta/2).$
If $\Delta\geq\pi$, the convex hull contains the origin, so $\rho(V)=0$. In either case,
$\sqrt{1-\rho(V)^{2}}=\sin(\Delta/2),$
with the convention $\sin(\Delta/2)=1$ for $\Delta\geq\pi$.
Substituting back into the Helstrom expression yields
$p^{*}(U_{0},U_{1})=\tfrac{1}{2}\big(1+\sin(\Delta/2)\big),$
as claimed. ∎
Witness definition.
Define the fixed-order ceiling
$p_{\rm fixed}^{*}\ :=\ \max\{\,p^{*}(I,AB),\ p^{*}(I,BA)\,\},$ (15)
and the switch performance
$p_{\rm switch}(\omega)\ :=\ p^{*}(I\otimes I,\ S(\theta(\omega))).$ (16)
The witness gap is
$\Delta(\omega)\ :=\ p_{\rm switch}(\omega)-p_{\rm fixed}^{*}.$ (17)
Proposition 6.2 (Convexity witness).
Let $\mathcal{F}$ be the set of fixed-order processes. For equal-prior discrimination,
$W\in{\rm conv}(\mathcal{F})\Rightarrow p^{*}(W)\leq\sup_{F\in\mathcal{F}}p^{*}(F)=p_{\rm fixed}^{*}.$
Hence if $\Delta(\omega)>0$, then $T(\omega)\notin{\rm conv}(\mathcal{F})$.
Proof.
For any discrimination strategy—specified by an input state and measurement—the success probability is given by the generalized Born rule $p(W)=\langle G,W\rangle$ for some Hermitian operator $G$ determined by the strategy. The optimal success probability
$p^{*}(W)=\sup_{G}\,\langle G,W\rangle$
is therefore the pointwise supremum of linear functionals of $W$, and hence convex. For a convex mixture $W=\sum_{i}q_{i}F_{i}$ with $F_{i}\in\mathcal{F}$ and $q_{i}\geq 0$, $\sum_{i}q_{i}=1$, one has
$p^{*}(W)\leq\sum_{i}q_{i}\,p^{*}(F_{i})\leq p_{\rm fixed}^{*}.$
Thus a positive witness gap $\Delta(\omega)>0$ implies that $T(\omega)$ cannot belong to $\mathrm{conv}(\mathcal{F})$. ∎
7 Main Results
We evaluate the Burau representation $\beta_{i}(s)$ of $B_{3}$, as discussed above, for $s=e^{i\omega/2}$ and verify $J(\omega)$–unitarity, positive-definiteness of $J(\omega)$, and Euclidean unitarity of the conjugated matrices $U(\omega)=R(\omega)\beta(w)R(\omega)^{-1}$ as defined in (10). The control braid word is $w=\sigma_{1}\sigma_{2}\sigma_{1}$, chosen as the minimal non-trivial representative of the Yang–Baxter class.
Single-qubit rotations.
Let us use the conventional $\mathrm{SU}(2)$ rotation matrices about the xx–, $y$--, and $z$-axes,
$R_{x}(\theta)$ $=\exp\!\left(-\tfrac{i}{2}\theta\,\sigma_{x}\right)\;=\;\begin{pmatrix}\cos\tfrac{\theta}{2}&-i\sin\tfrac{\theta}{2}\\[4.0pt] -i\sin\tfrac{\theta}{2}&\cos\tfrac{\theta}{2}\end{pmatrix}$, (18)
$R_{y}(\theta)$ $=\exp\!\left(-\tfrac{i}{2}\theta\,\sigma_{y}\right)=\begin{pmatrix}\cos\tfrac{\theta}{2}&-\sin\tfrac{\theta}{2}\\[4.0pt] \sin\tfrac{\theta}{2}&\cos\tfrac{\theta}{2}\end{pmatrix}$, (19)
$R_{z}(\phi)$ $=\exp\!\left(-\tfrac{i}{2}\phi\,\sigma_{z}\right)\;=\;\begin{pmatrix}e^{-i\phi/2}&0\\[4.0pt] 0&e^{i\phi/2}\end{pmatrix}$, (20)
where $\sigma_{x}$ and $\sigma_{z}$ are the Pauli matrices.
Das Gedankenexperiment.
Let us set the target unitaries as
$A=R_{x}(1.1),\qquad B=R_{z}(0.9),$
so that $A,B]\neq 0$ while both remain simple $SU(2)$ rotations. The switch phase is fixed as $\theta(\omega)=\omega$.
For each $\omega\in[0,2\pi]$, we compute:
  1. 1.
    JJ–unitarity errors $\|\beta_{i}^{\dagger}J\beta_{i}-J\|$ for $i=1,2$;
  2. 2.
    Euclidean unitarity error $\|U^{\dagger}U-I\|$ of the unitarized braid word;
  3. 3.
    the single-shot Helstrom success $p_{\rm switch}(\omega)$ for discriminating the switch process $S(\theta(\omega))$ from identity; and
  4. 4.
    the fixed-order ceiling $p_{\rm fixed}^{*}=\max\{p^{*}(I,AB),p^{*}(I,BA)\}$.
The code performing these computations, together with reproducible figures, is publicly available on GitHub [14].
Theorem 7.1.
The witness gap $\Delta(\omega)=p_{\rm switch}(\omega)-p_{\rm fixed}^{*}$ for the test device $T(\omega)$ with the above parameters is strictly positive for $|\omega|<\frac{2\pi}{3}$.
Indeed, the Squier form for $B_{2}$ is
$J_{3}(\omega)=\begin{pmatrix}2\cos\tfrac{\omega}{2}&-1\\[4.0pt] -1&2\cos\tfrac{\omega}{2}\end{pmatrix}$, (21)
which implies that $J_{3}(\omega)\succ 0$ for $|\omega|<\tfrac{2\pi}{3}$. The rest follows from numerical computations.
Figure 2: Left: verification of Squier’s $J$-unitarity for $\beta_{1}$, $\beta_{2}$. Center: Euclidean metric error of the unitarized word $U(\omega)$ within the positivity intervals of $J(\omega)$. Right: witness gap $\Delta(\omega)$ showing a positive region (non-convexity of causal order) where $J(\omega)\succ 0$.
Within the positive-definite window $|\omega|<2\pi/3$, the Squier generators satisfy $J$-unitarity to numerical precision better than $10^{-14}$, and the conjugated word $U(\omega)$ is Euclidean-unitary with error $\|U^{\dagger}U-I\|\!<\!10^{-14}$. The Helstrom success for the switch exceeds the fixed-order ceiling by a gap $\Delta_{\max}\approx 0.1796$ at $\omega\approx 1.754$, confirming causal non-separability.
Interpretation.
The control mixer $M(\omega)$ embodies the minimal non-Abelian structure of $B_{3}$: two non-commuting generators $\beta_{1}(\omega)$ and $\beta_{2}(\omega)$ acting unitarily on a two-dimensional control space. The existence of a positive witness gap $\Delta(\omega)>0$ demonstrates that the combined process cannot be expressed as a convex mixture of fixed-order operations.
In the language of anyon theory, this corresponds to an effective superposition of exchange orders producing a matrix holonomy rather than a scalar phase.
In a hypothetical realization, each braid word $w$ would correspond to an exchange of three distinguishable excitations, and $M(\omega)$ would represent the induced two-dimensional fusion subspace.
The non-commutativity of $\beta_{1}(\omega)$ and $\beta_{2}(\omega)$ ensures that successive exchanges act by non-commuting unitaries on this subspace, which is the defining feature of non-Abelian anyons.
Hence, the present construction provides a purely algebraic Gedankenexperiment demonstrating how non-Abelian exchange statistics could manifest as measurable order-sensitivity, independent of microscopic substrate.
Acknowledgments
A.K. was supported by the Wolfram Institute for Computational Foundations of Science, and by the John Templeton Foundation.
A.K. is grateful to the Institute of Theoretical Physics at ETH Zürich, and especially Juan Felipe Carrasquilla Álvarez, for hospitality during the research stay in Fall 2025.